OpenStudy (anonymous):

Can someone help me with this? thank you guys :) A restaurant serves coffee in copper mugs. A waiter fills a cup of mass 0.1 kg, initially at 20°C, with 0.2 kg of coffee initially at 70°C. What is the final temperature after the coffee and the cup attain thermal equilibrium? Assume that the coffee has the same specific heat capacity as water.

OpenStudy (anonymous):

I think you just average the temperature, since there is no heat exchange with the environment. \[\frac{(0.1kg \times 20C) + (0.2kg \times 50C)}{0.1kg + 0.2kg}\]

OpenStudy (anonymous):

thankyou :)))

OpenStudy (anonymous):

The boundaries of the system are not well defined, therefore the question is ambiguous. Including the whole restaurant in your system, the temperature will be the temperature in the restaurant. Including the whole universe, the answer will be 0K Regarding the cup and the coffee as one closed system, you should take into account the difference in heat capacity of copper and water. The answer you get from the heat balance: Tc Mc Cc + Tw Mw Cw = Te (Mc Cc + Mw Cw) Tc=temeprature cup °C Mc=mass of cup kg Cc=specific heat cupper J/kg/°C Tw,Mw,Cw=same for coffee Te=final temperature

OpenStudy (anonymous):

@harrybruning, @xtel wrote "Assume that the coffee has the same specific heat capacity as water".

OpenStudy (anonymous):

@harrybruning , @RafaelNeumann . my answer 71.85 degree celsius XD

OpenStudy (anonymous):

@harrybruning @RafaelNeumann thank you so much guys :)